# AI-native Memory: A Pathway from LLMs Towards AGI

### v3

@@ Abstract
Large language models (LLMs) have demonstrated the world with the sparks of artificial general intelligence (AGI). One opinion, especially from some startups working on LLMs, argues that an LLM with nearly unlimited context length can realize AGI. However, they might be too optimistic about the long-context capability of (existing) LLMs – (1) Recent literature has shown that their effective context length is significantly smaller than their claimed context length; and (2) Our reasoning-in-a-haystack experiments further demonstrate that simultaneously finding the relevant information from a long context and conducting (simple) reasoning is nearly impossible. In this paper, we envision a pathway from LLMs to AGI through the integration of memory. We believe that AGI should be a system where LLMs serve as core processors. In addition to raw data, the memory in this system would store a large number of important conclusions derived from reasoning processes. Compared with retrieval-augmented generation (RAG) that merely processing raw data, this approach not only connects semantically related information closer, but also simplifies complex inferences at the time of querying. As an intermediate stage, the memory will likely be in the form of natural language descriptions, which can be directly consumed by users too. Ultimately, every agent/person should have its own large personal model, a deep neural network model (thus AI-native) that parameterizes and compresses all types of memory, even the ones cannot be described by natural languages. Finally, we discuss the significant potential of AI-native memory as the transformative infrastructure for (proactive) engagement, personalization, distribution, and social in the AGI era, as well as the incurred privacy and security challenges with preliminary solutions.
@@ 1 Introduction
Acknowledging the necessity of Memory, we then discuss the forms of Memory and how to facilitate the interaction between Memory and LLM (e.g., loading the right data from “disk” to “RAM”). As an intermediate stage, the memory will likely be in the form of natural language descriptions. This is in line with many existing information extraction and knowledge discovery works and we will construct a “Memory Palace” for each agent/person. Ultimately, every agent/person should have its own large personal model (LPM), a deep neural network model (thus AI-native) that parameterizes and compresses all types of memory, even the ones cannot be described by natural languages. From this compression perspective, this LPM can be a LLM too. Finally, we discuss the significant potential of AI-native memory as the transformative infrastructure for AI-native (proactive) engagement, personalization, distribution, and social in the AGI era, as well as the incurred privacy and security challenges with preliminary solutions.
@@ 3 AGI should be a System with AI-Native Memory
What is AI-Native Memory? We believe the ultimate form of AI-Native Memory is a deep neural network model (thus AI-native) that parameterizes and compresses all types of memory, even the ones cannot be described by natural languages. In order to ensure the privacy of the Memory across different users who interacted with the same AGI agent, we argue that the best practice is to maintain one Memory model for each individual user. Therefore, we refer to this Memory model between the AGI agent and a particular user as the Large Personal Model (LPM) of this user. The LPM records, organizes, indexes, and arranges every detail about the individual, ultimately providing interfaces for users to directly access memories and for downstream applications (such as personalized generation, recommendations, etc.) to utilize useful, complete contexts. In a sense, the LPM acts as an upgraded “Retrieval-Augmented” role. Its superiority lies in the transformation of original data through extensive “reasoning” (i.e., organizing, indexing, etc.), rather than merely recording. Note that the LPM will evolve as the user interacts with LPM, creating a Data Flywheel.
@@ 5.2 Compared Methods
We have also included long-context LLMs’ results using Gemini Pro and GPT-4o. Note that the number of the tokens in the raw data have already exceeded the context length in these LLMs, so we only keep the most recent data. Because of such truncation, there are 5 (33%) Memory questions that can never be answered correctly by these baselines. For a fair comparison, we have excluded them when calculating the score for these two models.
@@ 5.3 Experiment Setup
Table 2: Pilot Experiment Results. The ratings are from 0 to 5 and averaged among 15 questions in each category. From left to right, the four categories largely follow an increasing order of difficulty levels, because they require more and more comprehensive and global understandings of the user. The only exceptions are those time-related Memory questions. Please note that in long-context LLM baselines, the raw data has been truncated so we have excluded 5 (33%) Memory questions that can never be answered correctly by these baselines; therefore, those two numbers with * are inflated. The best results are in bold and the second best ones are underlined.
@@
Long-context LLM-4o (L0) Raw Notes (128k) 2.7* 4.07 4.33 4.4 3.875
@@ 5.4 Results
In summary, RAG methods, by design and also verified in our experiments, are great at answering specific questions (i.e., local questions) with explicit mentions of entities and relations. Long-context LLMs have advantages in offering a macro understanding of the user and answering high-level questions (i.e., global questions). Our LPM is able to achieve (almost) the best performance on both local and global questions, compared with all these baselines.
@@ Appendix A An Example of Multi-Needle Reasoning-in-a-haystack
Appendix B Prompt Template
@@
Appendix C Evaluation Criteria
@@
Appendix D Detailed Experiment Results

### v4

@@ Abstract
Large language models (LLMs) have demonstrated the world with the sparks of artificial general intelligence (AGI). One opinion, especially from some startups working on LLMs, argues that an LLM with nearly unlimited context length can realize AGI. However, they might be too optimistic about the long-context capability of (existing) LLMs – (1) Recent literature has shown that their effective context length is significantly smaller than their claimed context length; and (2) Our reasoning-in-a-haystack experiments further demonstrate that simultaneously finding the relevant information from a long context and conducting (simple) reasoning is nearly impossible. In this paper, we envision a pathway from LLMs to AGI through the integration of memory. We believe that AGI should be a system where LLMs serve as core processors. In addition to raw data, the memory in this system would store a large number of important conclusions derived from reasoning processes. Compared with retrieval-augmented generation (RAG) that merely processing raw data, this approach not only connects semantically related information closer, but also simplifies complex inferences at the time of querying. As an intermediate stage, the memory will likely be in the form of natural language descriptions, which can be directly consumed by users too. Ultimately, every agent/person should have its own lifelong personal model, a deep neural network model (thus AI-native) that parameterizes and compresses all types of memory, even the ones cannot be described by natural languages. Finally, we discuss the significant potential of AI-native memory as the transformative infrastructure for (proactive) engagement, personalization, distribution, and social in the AGI era, as well as the incurred privacy and security challenges with preliminary solutions.
* Equal Contribution
@@ 1 Introduction
Acknowledging the necessity of Memory, we then discuss the forms of Memory and how to facilitate the interaction between Memory and LLM (e.g., loading the right data from “disk” to “RAM”). As an intermediate stage, the memory will likely be in the form of natural language descriptions. This is in line with many existing information extraction and knowledge discovery works and we will construct a “Memory Palace” for each agent/person. Ultimately, every agent/person should have its own lifelong personal model (LPM), a deep neural network model (thus AI-native) that parameterizes and compresses all types of memory, even the ones cannot be described by natural languages. From this compression perspective, this LPM can be a LLM too. Finally, we discuss the significant potential of AI-native memory as the transformative infrastructure for AI-native (proactive) engagement, personalization, distribution, and social in the AGI era, as well as the incurred privacy and security challenges with preliminary solutions.
@@ 3 AGI should be a System with AI-Native Memory
What is AI-Native Memory? We believe the ultimate form of AI-Native Memory is a deep neural network model (thus AI-native) that parameterizes and compresses all types of memory, even the ones cannot be described by natural languages. In order to ensure the privacy of the Memory across different users who interacted with the same AGI agent, we argue that the best practice is to maintain one Memory model for each individual user. Therefore, we refer to this Memory model between the AGI agent and a particular user as the Lifelong Personal Model (LPM) of this user. The LPM records, organizes, indexes, and arranges every detail about the individual, ultimately providing interfaces for users to directly access memories and for downstream applications (such as personalized generation, recommendations, etc.) to utilize useful, complete contexts. In a sense, the LPM acts as an upgraded “Retrieval-Augmented” role. Its superiority lies in the transformation of original data through extensive “reasoning” (i.e., organizing, indexing, etc.), rather than merely recording. Note that the LPM will evolve as the user interacts with LPM, creating a Data Flywheel.
@@ 5.2 Compared Methods
We have also included long-context LLMs’ results using Gemini 1.5 Pro and GPT-4o. The prompts used in these long-context LLMs are the same as the prompts in GraphRAG. Note that the number of the tokens in the raw data have already exceeded the context length in these LLMs, so we only keep the most recent data. Because of such truncation, there are 5 (33%) Memory questions that can never be answered correctly by these baselines. For a fair comparison, we have excluded them when calculating the score for these two models.
@@ 5.3 Experiment Setup
Table 2: Pilot Experiment Results. The ratings are from 0 to 5 and averaged among 15 questions in each category. From left to right, the four categories largely follow an increasing order of difficulty levels, because they require more and more comprehensive and global understandings of the user. The only exceptions are those time-related Memory questions. Please note that in long-context LLM baselines, the raw data has been truncated so we have excluded 5 (33%) Memory questions that can never be answered correctly by these baselines; therefore, those two numbers with * are inflated, e.g., 3.13.13.13.1* means $3.1\times 10/15=2.073.1 × 10 / 15 = 2.07$. The best results are in bold and the second best ones are underlined.
@@
Long-context LLM-4o (L0) Raw Notes (128k) 3.1* 3.4 4.0 3.2 3.425
@@ 5.4 Results
In summary, RAG methods, by design and also verified in our experiments, are great at answering specific questions (i.e., local questions) with explicit mentions of entities and relations. Long-context LLMs have advantages in offering a macro understanding of the user and answering high-level questions (i.e., global questions). Our LPM is able to achieve the best performance on both local and global questions, compared with all these baselines.
@@ Appendix A An Example of Multi-Needle Reasoning-in-a-haystack
Appendix B Prompt Template in Reasoning-in-a-haystack Experiments
@@
Appendix C Evaluation Criteria in Reasoning-in-a-haystack Experiments
@@
Appendix D Detailed Experiment Results of Reasoning-in-a-haystack